If Neptune originally formed at 20 AU and then migrated to its present distance of 30 AU, how much was Neptune's orbital speed reduced during the migration?
Answer #1

Intial distance of Neptune = R = 20 AU ; Present distance = R' = 30 AU

Let V be the initial orbital speed and V' be the final.

We know that, orbital speed is given by:

V = sqrt (GM/R)

V = sqrt (GM/20AU)

V' = sqrt (GM/30AU)

dividing V'/V

V'/V = [sqrt(GM) / sqrt (30 AU)] x [sqrt(20 AU) / sqrt(GM)]

V'/V = sqrt (20/30) = 0.816

V' = 0.816 V

Hence, during the migration the orbital speed of neptune got reduced by a factor of 0.816 of its original speed.